I don't know if this is a proportion or not but I don't understand:
1/3+1/5=x/30
Thanks

Answer by lwsshak3(11628)

I don't know if this is a proportion or not but I don't understand:
1/3+1/5=x/30
when working with equations, to eliminate fractions, multiply each term by the LCD, which in this case=30
30*1/3+30*1/5=30*x/30
10+6=x
x=16